# Defining subsystems in Hilbert spaces with non-Euclidean metric

### v2

@@ III Tomography and the no-signaling principle
Take the example of the spin 1/2 system. The set $\{\mathbb{1},\vec{\sigma}\}$ forms a tight frame in the trivial metric Hilbert space $\mathcal{H}=\mathbb{C}^{2}$. This is also the minimal group that induces an irreducible representation on $\mathcal{H}$[39]. The tight frame consists of measurement of the spin in three orthogonal directions since the Pauli matrices form the orthogonal basis for the space of observables in two dimensions: $tr(\sigma_{i}\sigma_{j})=2\delta_{ij}$. A general observable is therefore given by $\vec{\sigma}.\hat{n}$ and is associated with a direction $\hat{n}$ in 3-dimensional space. Let $p(m|\hat{n}_{i})$ be the probability of getting the outcome $m$, given the spin was measured in the $\hat{n}_{i}$ direction, corresponding to the measurement of the $\sigma_{i}$ observable. To make this measurement we use a Stern-Gerlach set-up where the particles of a specific velocity pass through the magnetic field $\vec{B}=B\hat{n}_{i}^{\perp}$ (see the appendix C for details) and then through a magnetic field in $z$-direction which splits the beam in two parts corresponding to the values of $m\in\{\pm\dfrac{1}{2}\}$. By controlling the direction and magnitude of the magnetic field $\vec{B}$, one produces the experimental data and constructs the density matrix: $\rho=\mathbb{1}/2+\sum_{m,i}mp(m|\hat{n}_{i})\sigma_{i}$. Note that while associating the operator $\sigma_{i}$ with the direction $\hat{n}_{i}$, the experimentalist makes two choices. The first is the choice of a tight frame from the family of unitarily equivalent tight frames. Since the set of operators $\{\mathbb{1},U\sigma_{i}U^{\dagger}\}$ also forms a tight frame in $\mathcal{H}$, the density matrix thus constructed is fixed only up to a unitary equivalence. The second choice is the metric space in which the operators $\{\sigma_{i}\}$ and their unitary equivalents are observables. One can equivalently choose the set of operators $\{\bar{\sigma}_{i}=\eta^{-1}\sigma_{i}\eta\}$ which are the observables in the metric space $\mathcal{H}_{G}$ with metric $G=\eta^{2}$. They also form a tight frame with $tr(\bar{\sigma}^{*}_{i}\bar{\sigma}_{j})=2\delta_{ij}$. With this choice of the metric space, the resultant matrix $\bar{\rho}=\mathbb{1}/2+\sum_{m,i}mp(m|n_{i})\bar{\sigma}_{i}$ is a density matrix in $\mathcal{H}_{G}$.
@@
Now consider a composite system of two distinguishable spin-1/2 particles. In $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ the set of operators $\{\mathbb{1},\sigma_{i}\otimes\sigma_{j},\mathbb{1}\otimes\sigma_{i},\sigma_{i}\otimes\mathbb{1}\}$ forms the tight frame. The experimental setup consists of two independent Stern-Gerlach apparatus mentioned above with independent control over the magnetic field $\vec{B}_{1}$ and $\vec{B}_{2}$. We then collect the data to construct the probabilities $p(m_{1},m_{2}|\hat{n}^{1}_{i},\hat{n}^{2}_{j})$ of obtaining the result $m_{1}$ as a result of measurement in $\hat{n}^{1}_{i}$ direction on particle 1 and $m_{2}$ as a result of measurement in $\hat{n}^{2}_{j}$ direction on particle 2. Born’s rule gives us the trace $\sum m_{1}m_{2}p(m_{1},m_{2}|\hat{n}^{1}_{i},\hat{n}^{2}_{j})=tr(\rho_{12}\sigma_{i}\otimes\sigma_{j})$. If the two particles are accessible to two distinct parties, namely Alice and Bob, then Alice only has the data $\{m_{1},p(m_{1}|\hat{n}^{1}_{i})\}$ which is a subset of the complete set of data of the system. When can we say that this subset of the complete data describes a subsystem? From the point of view of Alice, there has been no difference between the previous experiment and this experiment, and therefore she constructs a state in $\mathcal{H}_{1}$ namely $\rho_{1}$ using the Born rule $\sum m_{1}p(m_{1}|\hat{n}^{1}_{i})=tr(\rho_{1}\sigma_{i})$. From the global point of view, this expectation value is represented as $\sum m_{1}p(m_{1}|\hat{n}^{1}_{i})=tr(\rho_{12}\sigma_{i}\otimes\mathbb{1})$. ($\mathbb{1}$ corresponds to switching off the secondary magnetic field in the S-G set-up). Therefore, Alice’s statistical data corresponds to the set of observables $\{\sigma_{i}\otimes\mathbb{1}\}$, which forms a sub-algebra. With $\{\mathbb{1}\otimes\sigma_{i}\}$ forming the complementary sub-algebra, one can say that this data completely characterizes the sub-system, in this case, represented by the density matrix $\rho_{1}$.
@@
Consider the non-trivial metric representation of the system of two particles in $\mathcal{H}_{G}$ as the state $\bar{\rho}_{12}$. Alice, still in possession of the data $\{m,p(m,\hat{n}^{1}_{i})\}$, might be aware of the global metric structure $G$, but due to the non-separability of the metric, she does not have a notion of local observables. Therefore, she cannot construct a density matrix for particle 1 alone which can be seen as a reduced state of $\bar{\rho}_{12}$. Nevertheless, from the global point of view, we have $\sum m_{1}p(m_{1}|\hat{n}^{1}_{i})=tr(\bar{\rho}_{12}\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta)$. This is because we have $tr(\bar{\sigma}_{j}^{*}\bar{\sigma}_{j})=2\delta_{ij}$, which means that these measurements correspond to measurements in orthogonal directions. Therefore, Alice’s data also corresponds to the set of observables $\{\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta\}$, which forms a sub-algebra. With $\{\eta^{-1}(\mathbb{1}\otimes\sigma_{i})\eta\}$ forming the complementary sub-algebra together generating the total $C^{*}$ algebra $\mathcal{B}(\mathcal{H}_{G})$, Alice’s data corresponds to a well-defined subsystem of $\mathcal{H}_{G}$. But as we have established, the data $\{m,p(m|\hat{n}_{i}^{1})\}$ is faithfully represented by $\rho_{1}\in\mathcal{H}_{1}$, therefore, it captures a well-defined subsystem of $\mathcal{H}_{G}$. These arguments can be generalized for an arbitrary finite-dimensional system, and therefore, we have the following theorem:
@@
Given a density matrix $\bar{\rho}_{12}$ in $\mathcal{H}_{G}$, the density matrix $\rho_{1}=tr_{2}(\eta\bar{\rho}_{12}\eta^{-1}$) represents a well-defined sub-system of the system represented by $\bar{\rho}_{12}$.
The no-signaling principle: The no-signaling principle states that the statistics of the subsystem measured by Alice denoted by $\{m_{1},p(m_{1}|\hat{n}^{1}_{i})\}$ should not change by an operation on another subsystem belonging to Bob. One can show that this principle holds in all metric spaces. If Alice’s and Bob’s subsystems correspond to the sub-algebra $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ respectively, then for their shared system $\bar{\rho}_{12}$ in the metric space $(\mathcal{H}_{G},\pi_{G})$, such that $\eta\pi_{G}(B)\eta^{-1}=\mathbb{1}\otimes\mathcal{O}_{B}$, we have the following relation:
$\rho_{1}=tr_{2}(\eta\bar{\rho}_{12}\eta^{-1})=tr_{2}(\eta\pi_{G}(B)\bar{\rho}_{12}\eta^{-1}),\,\forall\,B\in\mathcal{A}_{2}.$ (4)
@@
The second equality comes from the no-signaling principle in trivial metric space: $tr_{2}(\eta\pi_{G}(B)\bar{\rho}_{12}\eta^{-1})=tr_{2}(\mathbb{1}\otimes\mathcal{O}_{B}\rho_{12})=tr_{2}(\rho_{12})$. Since we have already established that the state $\rho_{1}$ faithfully represents Alice’s subsystem, we can conclude that the no-signaling principle for a bi-partition, as given in the definition 1, is obeyed in every metric space. In the literature [15], one typically shows that the no-signaling principle is satisfied when the metric is separable $G=G_{1}\otimes G_{2}$, which is only a specific case of the more general principle that we have exhibited here through equation (4).
IV Conclusion
@@
The axiom of the tensor product decomposition of the Hilbert space of composite quantum systems restricts how we understand a subsystem. In a vast family of quantum systems, such a decomposition is not possible, and consequently, a partial trace operation is not well defined. This issue arises in the non-trivial metric space representation of composite systems, which forms a natural description of the systems undergoing balanced gain and loss under the pseudo-Hermitian Hamiltonian. In this work, we have used algebraic tools to provide an operational definition of subsystems in such Hilbert spaces with non-trivial metric structures. We show that given an algebra and a state over it, one can equivalently construct a Hermitian or a pseudo-Hermitian cyclic representation, which are unitarily equivalent if we use the term unitary in its most general sense. While given a pseudo-Hermitian Hamiltonian, one can choose from a family of Hamiltonian compatible metric spaces, we show that different choices of the metric operator can correspond to choosing different system decompositions. Since choosing a metric is equivalent to choosing a reference frame for the system, it makes sense that choosing two different metric representations changes the subs-system properties. Furthermore, the choice of sub-system decomposition is equivalent to a choice of an equivalence class of metric spaces. Finally, we show that from an experimentalist’s point of view, local tomography on a sub-system undergoing pseudo-Hermitian evolution creates no paradoxes, and allows one to construct a local description of the subsystem in trivial metric space. The no-signaling principle, which so far has been restricted to tensor product decomposable Hilbert spaces, can be consistently shown to hold good for an algebraically defined sub-system, even when the Hilbert spaces do not have a genuine tensor product structure.
@@
After the mathematical description of pseudo-Hermitian quantum mechanics as a trivial extension of normal quantum mechanics was established, most studies dealing with composite systems in this field have been wary of using the non-separable metric structure due to the ill-defined partial trace operation in such scenarios. From this work, it is clear that such metric operators are as physically meaningful as the separable ones, and there is no clear physical reasoning for choosing a separable metric operator over the non-separable operators to define the subsystems. We hope that this work will be helpful in a better understanding of composite quantum systems evolving under the $\mathcal{P}\mathcal{T}$-symmetric non-Hermitian Hamiltonians.
@@ Appendix A Lindbladian dynamics in $\mathcal{H}_{G}$
define $H_{e}=H-i\sum_{i}L_{j}^{\dagger}L_{j}$ (effective non-Hermitian Hamiltonian) so that the above equation is can be recast in the following form
@@
The probability of the jump in this infinitesimal time interval $t$ to $t+dt$ is given by $dp=\sum_{j}tr(L_{j}^{\dagger}L\rho(t))dt$ and the probability of no jump in this interval is $p_{0}=1-dp=tr(M_{0}^{\dagger}M_{0}\rho(t))$.
If the effective Hamiltonian is pseudo-Hermitian ($H_{e}^{\dagger}\eta^{2}=\eta^{2}H_{e}$), then, the transformation $\rho\rightarrow\rho\eta^{2}=\bar{\rho}$ rewrites the Lindblad equation as
@@
Now the jump probability within the time interval $t$ to $t+dt$ is $dp=\sum_{j}tr(M_{j}^{*}M_{j}\bar{\rho})=\sum_{j}tr(L_{j}^{\dagger}\eta^{2}L_{j}\rho)$ and the probability of no jump is $p_{0}=tr(M_{0}^{*}M_{0}\bar{\rho})$. However, one can check that $M_{0}^{*}M_{0}=\mathbb{1}$, therefore, $p_{0}=tr(\bar{\rho})$. Since one can always choose $\eta^{2}$ s.t. $tr(\bar{\rho})=1$, the pseudo-Hermiticity condition tells us $\sum_{j}tr(L_{j}^{\dagger}L_{j}\eta^{2})=0$. From here we can conclude $\sum_{j}tr(L_{j}^{\dagger}\eta^{2}L_{j}\rho)=0$ In other words, the jump probability in the redefined Lindbladian equation vanishes if the effective Hamiltonian is Hermitian, indicating a unitary evolution.
Appendix B Ambiguities in entropy of the state
Given a finite-dimensional representation of the C*-algebra $\mathcal{A}$, given by $(\mathcal{H},\pi)$, a state $\omega$ on this algebra is uniquely represented by the element $\rho_{\omega}\in\pi(\mathcal{A})$ of this algebra, such that
$\omega(A)=tr(\rho_{\omega}\pi(A))\hskip 28.45274pt\forall\ A\in\mathcal{A}.$
Now, if the state $\omega$ is pure, so is the density matrix $\rho_{\omega}$. If it is a mixed state, it can be written as a convex combination of other pure states: $\omega=\sum_{\alpha}\lambda_{\alpha}\omega^{\alpha}$, and the corresponding density matrix can be written as $\rho_{\omega}=\sum_{\alpha}\lambda_{\alpha}{\rho_{\omega}}^{\alpha}$ where $\rho_{\omega}^{\alpha}$. are pure state density matrices Furthermore, the representation is reducible into a direct sum of irreducible representations $\rho_{\omega}=\oplus_{i}p_{i}\rho_{i}$. Note that here $\rho_{i}$ need not be pure states. The convex sum decomposition of the state $\omega$ is not unique since the state space is not a simplex. Therefore, a state can be prepared from an ensemble of pure states in many ways, represented by the probability vector $\vec{\lambda}$ s.t. $\sum_{\alpha}\lambda_{\alpha}=1$. The entropy associated with such a preparation is given by the Shannon entropy of the probability vector $\vec{\lambda}$: $H(\vec{\lambda})=-\sum_{\alpha}\lambda_{\alpha}\ln(\lambda_{\alpha})$ The von-Newmann entropy of the density matrix $\rho_{\omega}$, on the other hand, is given by $S(\rho_{\omega})=-tr(\rho_{\omega}\ln\rho_{\omega})=\sum_{i}e_{i}\ln e_{i}$ where $\vec{e}$ is the spectrum of the state $\rho_{\omega}$. A density matrix can be written as a convex combination of pure density matrices in many non-unique ways: $\rho=\sum_{i}\lambda_{i}|\psi_{i}\rangle\langle\psi_{i}|$. It turns out that the Shannon entropy of all such preparations (given by $H(\vec{\lambda})$ ) is bounded below by the von-Newmann entropy of the state: $S(\rho)=\inf\{H(\vec{\lambda});\rho=\lambda_{i}|\psi_{i}\rangle\langle\psi_{i}|\}$. Similarly one can construct the unique entropy of the state $\omega$ on the algebra $\mathcal{A}$ as the infimum of all the preparations of $\omega$: $S(\omega)=\inf\{H(\vec{\lambda});\omega=\sum_{\alpha}\lambda_{\alpha}\omega^{\alpha}\}$. We therefore have a unique notion of the entropy of a state if $S(\omega)=S(\rho_{\omega})$. As shown in [34], this is true only under faithful representations with no multiplicities. The most general relation between the two entropies is given as follows:
$S(\rho_{\omega})=S(\omega)+\sum_{i}p_{i}\ln m_{i}$ (10)
where $m_{i}$ are the multiplicities of the $i-th$ irreducible subspace and $p_{i}$ are the weights of each invariant subspace in the representation $\rho_{\omega}=\oplus_{i}p_{i}\rho_{i}$. In this work, we will define the entropy of the state $\omega$ as the von Neumann entropy of the unique density matrix $\rho_{\omega}$, implying that we will restrict to the multiplicity-free faithful representation.

### v3

@@ IV.1 An example
Note that while associating the operator $\sigma_{i}$ with the direction $\hat{n}_{i}$, the experimentalist makes two choices. The first is the choice of a tight frame from the family of unitarily equivalent tight frames. Since the set of operators $\{\mathbb{1},U\sigma_{i}U^{\dagger}\}$ also forms a tight frame in $\mathcal{H}$, the density matrix thus constructed is fixed only up to a unitary equivalence. The second choice is that of the metric space in which the operators $\{\sigma_{i}\}$ and their unitary equivalents are observables (in this case it is the Euclidean metric space). Equivalently, one can choose the set of operators $\{\bar{\sigma}_{i}=\eta^{-1}\sigma_{i}\eta\}$ which are the observables in the metric space $\mathcal{H}_{G}$ with metric $G=\eta^{2}$. The set $\{\bar{\sigma}_{i}\}$ also forms a tight frame and their measurements can be associated with measurements in orthogonal directions since these operators are orthogonal with respect to the Hilbert-Schmidt inner product on $\mathcal{H}_{G}$: $\operatorname{tr}(\bar{\sigma}^{*}_{i}\bar{\sigma}_{j})=2\delta_{ij}$. With this choice of the metric space, the resultant operator $\bar{\rho}=\mathbb{1}/2+\sum_{m,i}mp(m|n_{i})\bar{\sigma}_{i}$ is a density matrix in $\mathcal{H}_{G}$.
@@
Now, consider a composite system of two distinguishable spin-1/2 particles. In the Euclidean metric space $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ the set of operators $\mathcal{A}_{12}:=\{\mathbb{1}_{12},\sigma_{i}\otimes\sigma_{j},\mathbb{1}_{1}\otimes\sigma_{i},\sigma_{i}\otimes\mathbb{1}_{2}\}$ forms the tight frame. The experimental setup consists of two independent Stern-Gerlach apparatus mentioned above. We then collect the data to construct the probabilities $p(m_{1},m_{2}|\hat{n}^{1}_{i},\hat{n}^{2}_{j})$ of obtaining the result $m_{1}$ as a result of measurement in $\hat{n}^{1}_{i}$ direction on particle 1 and $m_{2}$ as a result of measurement in $\hat{n}^{2}_{j}$ direction on particle 2. If the two particles are accessible to two distinct parties, namely Alice and Bob (whose system indices will be labeled as $1$ and $2$ respectively), then Alice only has the data $\{m_{1},p(m_{1}|\hat{n}^{1}_{i})\}$ which is a subset of the complete set of data of the system. Can we say that this subset of the complete data describes a subsystem? Alice can construct a state in $\mathcal{H}_{1}$ namely $\rho_{1}$ using the Born rule $\sum_{m_{1}}m_{1}p(m_{1}|\hat{n}^{1}_{i})=\operatorname{tr}(\rho_{1}\sigma_{i})$. From the global point of view, this expectation value is represented as $\sum_{m_{1}}m_{1}p(m_{1}|\hat{n}^{1}_{i})=\operatorname{tr}(\rho_{12}\sigma_{i}\otimes\mathbb{1})$. Therefore, Alice’s data corresponds to the set of observables $\{\sigma_{i}\otimes\mathbb{1}\}$, which forms a subalgebra of the tight frame $\mathcal{A}_{12}$. With the set $\{\mathbb{1}\otimes\sigma_{i}\}$, corresponding to the data $\{m_{2},p(m_{2}|\hat{n}_{j}^{2})\}$ belonging to Bob, forming the complementary subalgebra of $\mathcal{A}_{12}$, one can say that this data completely characterizes Alice’s subsystem, in this case, represented by the density matrix $\rho_{1}$.
@@
Consider the non-Euclidean metric representation of the system of two particles in $\mathcal{H}_{G}\neq\mathcal{H}_{G_{1}}\otimes\mathcal{H}_{G_{2}}$, as the state $\bar{\rho}_{12}$ related to the density matrix defined above through $\bar{\rho}_{12}=\eta^{-1}\rho_{12}\eta$. Alice, still in possession of the data $\{m,p(m,\hat{n}^{1}_{i})\}$, might be aware of the global metric structure $G$, but due to the structure of the metric, she does not have a notion of local observables. Therefore, she cannot construct a density matrix for particle 1 alone which can be seen as a reduced state of $\bar{\rho}_{12}$. Nevertheless, from the global point of view, her data can be given as
$\sum_{m_{1}}m_{1}p(m_{1}|\hat{n}^{1}_{i})=\dfrac{1}{2}\operatorname{tr}(\bar{\rho}_{12}\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta).$ (18)
Therefore, Alice’s data also corresponds to the set of observables $\{\eta^{-1}(\sigma_{i}\otimes\mathbb{1})\eta\}$, which forms a subalgebra of the total algebra $\eta^{-1}\mathcal{A}\eta$. With $\{\eta^{-1}(\mathbb{1}\otimes\sigma_{i})\eta\}$ forming the complementary subalgebra together generating the total $C^{*}$ algebra $\mathcal{B}(\mathcal{H}_{G})$, Alice’s data corresponds to a well-defined subsystem of $\mathcal{H}_{G}$. But as we have established, the data $\{m,p(m|\hat{n}_{i}^{1})\}$ is faithfully represented by $\rho_{1}=\operatorname{tr}_{2}(\rho_{12})\in\mathcal{H}_{1}$. These arguments can be generalized for an arbitrary finite-dimensional system, and therefore, we have the following theorem:
@@
Given a density matrix $\bar{\rho}_{12}$ in $\mathcal{H}_{G}$, the density matrix $\rho_{1}=\operatorname{tr}_{2}(\eta\bar{\rho}_{12}\eta^{-1}$) represents a well-defined subsystem of the system represented by $\bar{\rho}_{12}$ if $\eta\bar{\rho}\eta^{-1}\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$.
As we discussed, the Hermitisization of the state to a Hermitian density matrix corresponds to a unitarily equivalent representation of the state. Therefore, the bipartite states $\rho_{12}$ and $\bar{\rho}_{12}$ have the same entanglement structure. For instance, if the state $\bar{\rho}_{12}$ (or equivalently the state $\rho_{12}$) is a pure state, the entropy of the reduced state $\rho_{1}$ is a measure of entanglement between the subsystems 1 and 2 in both representations.
IV.2 The no-signalling principle
Suppose there is a constraint on the composite system shared between Alice and Bob, that Alice cannot signal to Bob. This can happen due to the constraints of the experimental setup, e.g. if Alice and Bob’s measurement events are space-like separated, or if Alice is in the causal future of Bob. In such a scenario, the no-signalling principle states that the statistics of the subsystem measured by Alice, denoted by $\{m_{1},p(m_{1}|\hat{n}^{1}_{i})\}$, should not change by a measurement on another subsystem belonging to Bob. Taking the example of the composite spin-1/2 system that we just discussed, no-signalling from Alice to Bob means that the conditional probabilities $p(m_{1}|\hat{n}^{1}_{i})$ must be independent of Bob’s measurement directions $\hat{n}_{j}^{2}$:
$\sum_{m_{2}}p(m_{1},m_{2}|\hat{n}^{1}_{i},\hat{n}^{2}_{j})=p(m_{1}|\hat{n}^{1}_{i})$ (19)
Suppose Alice’s and Bob’s subsystems correspond to the subalgebra $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ respectively. The POVM elements $\{M_{j}^{m}\}$ corresponding to the measurements by Bob in the $\hat{n}_{j}$ direction yielding the result $m$, belong to the subalgebra $\mathcal{A}_{2}$. Given a representation of the total system in $(\mathcal{H}_{G},\pi_{G})$, the POVM elements of Bob given by $\pi_{G}(M_{j}^{m})$ satisfy the relation $\eta\pi_{G}(M_{j}^{m})\eta^{-1}=\mathbb{1}_{1}\otimes\Pi_{j}^{m}$, where $\Pi_{j}^{m}$ are the POVMs elements of Bob as represented as operators on $\mathcal{H}_{2}$, such that $\sum_{m}\Pi_{j}^{m}=\mathbb{1}_{2}$. Therefore, for their shared system $\bar{\rho}_{12}$ in the metric space $(\mathcal{H}_{G},\pi_{G})$, we have the following relation:
$\rho_{1}=\operatorname{tr}_{2}(\eta\bar{\rho}_{12}\eta^{-1})$ $=\sum_{m}\operatorname{tr}_{2}(\eta\pi_{G}(M_{j}^{m})\bar{\rho}_{12}\eta^{-1})$ (20)
The second equality comes from the no-signalling principle in Euclidean metric space stated as the following :
$\sum_{m}\operatorname{tr}_{2}(\eta\pi_{G}(M_{i}^{m})\bar{\rho}_{12}\eta^{-1})=\sum_{m}\operatorname{tr}_{2}(\mathbb{1}\otimes\Pi_{i}^{m}\rho_{12})=\operatorname{tr}_{2}(\rho_{12})$ (21)
@@ IV.2 The no-signalling principle
Since we have already established that the state $\rho_{1}$ faithfully represents Alice’s subsystem, we can conclude that the no-signalling principle for a bi-partition, as given in the definition 3, is obeyed in every metric space. In the literature, one typically shows that the no-signalling principle is satisfied when the metric is separable $G=G_{1}\otimes G_{2}$, e.g. [15], which is only a specific case of the more general principle that we have exhibited here through equation (20).
V Conclusion
@@
The axiom of the tensor product decomposition of the Hilbert space of composite quantum systems restricts how we understand a subsystem. In a vast family of quantum systems, such a decomposition is not possible, and consequently, a partial trace operation is not well defined. This issue arises in the non-trivial metric space representation of composite systems, which forms a natural description of the systems undergoing balanced gain and loss under the pseudo-Hermitian Hamiltonian. In this work, we have used algebraic tools to provide an operational definition of subsystems in such Hilbert spaces with non-trivial metric structures. We show that given an algebra and a state over it, one can equivalently construct a Hermitian or a pseudo-Hermitian cyclic representation, which are unitarily equivalent if we use the term unitary in its most general sense. While, given a pseudo-Hermitian Hamiltonian, one can choose from a family of Hamiltonian compatible metric spaces, we show that different choices of the metric operator can correspond to choosing different system decompositions. Since choosing a metric is equivalent to choosing a reference frame for the system, it makes sense that choosing two different metric representations changes the subsystem properties. Furthermore, the choice of subsystem decomposition is equivalent to a choice of an equivalence class of metric spaces. Finally, we show that from an experimentalist’s point of view, local tomography on a subsystem undergoing pseudo-Hermitian evolution creates no paradoxes, and allows one to construct a local description of the subsystem in trivial metric space. The no-signalling principle, which so far has been restricted to tensor product decomposable Hilbert spaces, can be shown to hold good for an algebraically defined subsystem, even when the Hilbert spaces do not have a genuine tensor product structure.
@@
After the mathematical description of pseudo-Hermitian quantum mechanics as a trivial extension of normal quantum mechanics was established, most studies dealing with composite systems in this field have been wary of using the non-separable metric structure due to the ill-defined partial trace operation in such scenarios. From this work, it is clear that such metric operators are as physically meaningful as the separable ones, and there is no clear physical reasoning for choosing a separable metric operator over the non-separable operators to define the subsystems. We hope that this work will be helpful towards a better understanding of composite quantum systems evolving under the $\mathcal{P}\mathcal{T}$-symmetric non-Hermitian Hamiltonians.
@@ Appendix A Lindbladian dynamics in $\mathcal{H}_{G}$
define $H_{e}=H-\dfrac{i}{2}\sum_{i}L_{j}^{\dagger}L_{j}$ (effective non-Hermitian Hamiltonian) so that the above equation can be recast in the following form
@@
The probability of the jump in this infinitesimal time interval $t$ to $t+dt$ is given by $dp=\sum_{j}\operatorname{tr}(L_{j}^{\dagger}L_{j}\rho(t))dt$ and the probability of no jump in this interval is $p_{0}=1-dp=\operatorname{tr}(M_{0}^{\dagger}M_{0}\rho(t))$.
If the effective Hamiltonian is quasi-Hermitian ($H_{e}^{\dagger}\eta^{2}=\eta^{2}H_{e}$), then, the transformation $\rho\rightarrow\rho\eta^{2}=\bar{\rho}$ rewrites the Lindblad equation as
@@
Now the jump probability within the time interval $t$ to $t+dt$ is $dp=\sum_{j}\operatorname{tr}(M_{j}^{*}M_{j}\bar{\rho})=\sum_{j}\operatorname{tr}(L_{j}^{\dagger}\eta^{2}L_{j}\rho)dt$ and the probability of no jump is $p_{0}=\operatorname{tr}(M_{0}^{*}M_{0}\bar{\rho})$. However, one can check that $M_{0}^{*}M_{0}=\mathbb{1}$, therefore, $p_{0}=\operatorname{tr}(\bar{\rho})$. Since one can always choose $\eta^{2}$ s.t. $\operatorname{tr}(\bar{\rho})=1$, the quasi-Hermiticity condition tells us $\sum_{j}\operatorname{tr}(L_{j}^{\dagger}L_{j}\eta^{2})=0$. From here we can conclude $\sum_{j}\operatorname{tr}(L_{j}^{\dagger}\eta^{2}L_{j}\rho)=0$. In other words, the jump probability in the redefined Lindbladian equation vanishes if the effective Hamiltonian is Hermitian, indicating a unitary evolution.
Appendix B GNS constrcution
The Gelfand–Naimark–Segal or GNS construction  [36], is the construction of a Hilbert space description of this system, which involves first constructing a $C^{*}$ algebra on which the state is faithful, i.e., has a trivial kernel. This algebra $\mathcal{U}_{\omega}$ is constructed by first identifying the kernel of the state
$K_{\omega}=\{X\in\mathcal{A}\,\text{ s.t. }\,\omega(X^{*}X)=0\},$ (27)
and “quotienting out” the kernel from the algebra: $\mathcal{U}_{\omega}:=\mathcal{A}/K_{\omega}$. In other words, the elements of $\mathcal{U}_{\omega}$ consist of the equivalence classes
$\psi_{A}=\{A+X\,\text{ s.t. }\,X\in K_{\omega}\}$ (28)
One can see that the functional $\omega$ is faithful on $\mathcal{U}_{\omega}$. Moreover, $\mathcal{U}_{\omega}$ is a complex vector space, a property one can check by using the fact that $AX\in K_{\omega}\,\forall\,X\in K_{\omega}$, i.e. $K_{\omega}$ is a left ideal of $\mathcal{A}$. In the next step of the GNS construction, the state functional $\omega$ is used to define an inner product on the vector space $\mathcal{U}_{\omega}$ via $\langle\psi_{A}|\psi_{B}\rangle=\omega(A^{*}B)$. With this added structure, $\mathcal{U}_{\omega}$ becomes a pre-Hilbert space which can be completed to form a Hilbert space $\mathcal{H}_{\omega}=\overline{\mathcal{U}}_{\omega}$. The bar over the algebra denotes the completion with respect to the norm induced by the inner product. Therefore, the set of vectors $\{\psi_{A}\}$ is dense in $\mathcal{H}_{\omega}$. If one defines a representation given by the following action on the states: $\pi_{\omega}(A)\psi_{B}=\psi_{AB}$, one can then derive the relation
$\omega(A^{*}B)=\langle\psi_{\mathbb{1}}|\pi_{\omega}(A^{*}B)\psi_{\mathbb{1}}\rangle$ (29)
where $\mathbb{1}$ is the identity element of the algebra $\mathcal{A}$. Therefore, $\psi_{\mathbb{1}}$ is a cyclic vector of the representation $(\mathcal{H}_{\omega},\pi_{\omega})$. This construction gives us a faithful cyclic representation of the algebra $\mathcal{U}_{\omega}$, with the representation space $\mathcal{H}_{\omega}$, the map $\pi_{\omega}$, and the cyclic vector $\psi_{\mathbb{1}}$, such that $\omega(A)=\langle\psi_{\mathbb{1}}|\pi(A)\psi_{\mathbb{1}}\rangle_{\mathcal{H_{\omega}}}$.
B.1 Combining the two representations
Given two systems $\{\mathcal{A}_{1},\omega_{1}\}$ and $\{\mathcal{A}_{2},\omega_{2}\}$ and their respective representations $\{\mathcal{H}_{\omega_{1}},\pi^{1}_{\omega_{1}}\}$ and $\{\mathcal{H}_{\omega_{2}},\pi^{2}_{\omega_{2}}\}$, we will construct a larger representations space. We define the Kernel space of these states as $K_{\omega_{1}}=\{X\in\mathcal{A}_{1}\,\text{and}\,\omega(X^{*}X)=0\}$ and similarly $K_{\omega_{2}}$, and use them to construct the Hilbert spaces $\mathcal{H}_{\omega_{1}}=\overline{\mathcal{A}_{1}/K_{\omega_{1}}}$ and $\mathcal{H}_{\omega_{2}}=\overline{\mathcal{A}_{2}/K_{\omega_{2}}}$. Define the product of these spaces $\mathcal{H}_{\omega_{1}}\odot\mathcal{H}_{\omega_{2}}$ as a linear combination of the elements of the form $\psi_{A_{i}}^{1}\psi_{B_{j}}^{2}$[38] where we have $A_{i}\in\mathcal{A}_{1}$ and $B_{j}\in\mathcal{A}_{2}$. This forms a linear vector space with an inner product defined through the state $\omega(A_{i}B_{j})=\omega_{1}(A_{i})\omega_{2}(B_{j})$. Note that this is not a faithful state over $\mathcal{H}_{\omega_{1}}\odot\mathcal{H}_{\omega_{2}}$. This is because the kernel of $\omega$ cannot be constructed solely from the kernels $K_{\omega_{1}}$ and $K_{\omega_{2}}$. One, therefore, defines a Hilbert space called the direct product Hilbert space by quotienting out this kernel: $\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}=(\overline{\mathcal{H}_{\omega_{1}}\odot\mathcal{H}_{\omega_{2}})/K_{\omega}}$. Under this equivalence, the elements of the form $\psi_{A_{i}}^{1}\psi_{B_{j}}^{2}$ are mapped to the elements of the form $\psi_{A_{i}}^{1}\otimes\psi_{B_{j}}^{2}\in\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$.